2 balls have a mass of 0.838kg each. The gravitational force between them is 8.08e-11N. The universal constant G = 6.673e-11 Nm^2/kg^2. How far apart are the balls?
Answer #1

we know

gravitational force F = G*M*m/R2

R2 = G*M*m/F

R2 = 6.67*10-11*0.838*0.838/8.08*10-11

R = 0.7613 m

R = 76.13 cm